What is the Yield to Maturity (YTM) on a 3.65% T-Bond with a maturity date in October 2021 if the current market price is 111.82? How should you interpret this YTM?

A) YTM = 0.46%, the maximum premium over face value of the bond that an investor should pay.

B) YTM = 1.84%, the average annual return the investor will earn if the bond is purchased at the quoted price and held to maturity.

C) YTM = 0.92%, the average annual return the investor will earn if the bond is purchased at the quoted price and held to maturity.

D) YTM = 0.92%, the maximum premium over face value of the bond that an investor should pay.

is there a formula for YTM needed for this question?
Answer #1

Option B. YTM = 1.84%, the average annual return the investor will earn if the bond is purchased at the quoted price and held to maturity.

YTM is calculated as follows:

Assuming

Number of years to maturity = 7 and the bond is paying semi-annual coupon payment.

nper = 7 *2 = 14

pmt = 3.65 * 100 / 2 = 1.825

pv = -111.82

fv = 100 ( par value)

Semi-annual YTM is calculated in excel as follows:

=rate (nper,pmt,pv,fv)

=rate(14,1.825,-111.82,100)

=0.92

Annual YTM = 0.92 * 2 = 1.84%

Considering that the number of years to maturity is 7 years, then annual YTM would be 1.84%. Yield to maturity is the annual return expected on a bond when held till maturity.